# Robust Watermarking Using Generative Priors Against Image Editing: From Benchmarking to Advances

### v1

@@ Acknowledgments
This research is supported by the National Research Foundation, Singapore under its Strategic Capability Research Centres Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.
@@ A.2 Generative model-based image editing
Recent and significant advancements in text-to-image generative models (Chang et al., 2023; Ding et al., 2022; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Yu et al., 2022; Peebles & Xie, 2023; Esser et al., 2024) have enhanced numerous applications (Avrahami et al., 2023; Ruiz et al., 2023; Hertz et al., 2022; Kim et al., 2022; Tumanyan et al., 2023; Zhang et al., 2023; Mou et al., 2023; Shi et al., 2024; Zhu et al., 2023; Zhou et al., 2023; 2024a; Lu et al., 2023; 2024; Zhao et al., 2023a; 2024a; 2024b; Wang & Chen, 2024; Gandikota et al., 2023; Parmar et al., 2024; Zhou et al., 2024b; c; Zeng et al., 2024). In this study, we focus on real image editing, which allows users to freely modify actual photographs, producing highly realistic results. Typically, the inputs for image editing include an image and various conditions that help users accurately describe their desired changes. These conditions can encompass text prompts using natural language to specify the edits (Brooks et al., 2023; Zhang et al., 2024c; Fu et al., 2023; Zhang et al., 2024b), region masks to designate areas for modification (Zhao et al., 2024c; Zhuang et al., 2023; Wang et al., 2023), additional images to provide desired styles or objects (Chen et al., 2024; Lu et al., 2023; Yang et al., 2023), and drag points (Pan et al., 2023; Shi et al., 2024; Mou et al., 2023) that enable users to interactively move specific points in the image to target positions.
@@ Appendix B Watermarking Frequency Patterns
Figure 6: Frequency pattern visualizations for each watermarking method. The DWTDCT method is excluded because it closely resembles DWTDCTSVD and their pattern intensity is too weak to be discerned on the uniform scale. Please zoom in for a closer look.
@@
Appendix C Resolution Scaling
C.1 Method
@@
C.2 Impact on watermarked image quality
@@
C.3 Impact on robustness against distortions
@@
Appendix D Condition Adapter Architecture
@@
Appendix E Implementation Details
E.1 Pretraining
@@ E.2 Fine-tuning
Appendix F Analysis of Image Editing
F.1 Image regeneration
@@ F.2 Global editing and local editing
F.3 Image-to-video generation
@@
Appendix G Inference Speed and GPU Memory Evaluation
@@
Appendix H Qualitative Comparison
@@
Figure 15: Qualitative comparison of the evaluated watermarking methods. Please zoom in for a closer look.
Appendix I Additional Benchmarking Results

### v2

@@ Acknowledgments
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
@@ A.1 In-generation image watermarking
Figure 6 shows the results of applying in-generation watermarking methods combined with image inversion techniques to embed watermarks in real images. Specifically, we applied DDIM/DPM inversion methods to extract initial noises from real images, added watermarks using three in-generation watermarking techniques—Tree Ring (Wen et al., 2023), RingID (Ci et al., 2024), and Gaussian Shading (Yang et al., 2024)—and then inverted the noise back to produce the watermarked image. The resulting image differed significantly from the original, thereby undermining the photographer’s intent to protect their work.
Figure 6: Comparison of Tree Ring (Wen et al., 2023), RingID (Ci et al., 2024), and Gaussian Shading (Yang et al., 2024) methods for embedding watermarks in real images. Each real image is first inverted into noise using a 50-step image inversion process before the watermark is injected.
@@ A.2 Generative model-based image editing
Recent and significant advancements in text-to-image generative models (Chang et al., 2023; Ding et al., 2022; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022; Yu et al., 2022; Peebles & Xie, 2023; Esser et al., 2024) have enhanced numerous applications (Avrahami et al., 2023; Ruiz et al., 2023; Hertz et al., 2022; Kim et al., 2022; Tumanyan et al., 2023; Zhang et al., 2023; Mou et al., 2023; Shi et al., 2024; Zhu et al., 2023; Zhou et al., 2023; 2024a; Lu et al., 2023; 2024; Zhao et al., 2023a; 2024a; 2024b; Wang & Chen, 2024; Gandikota et al., 2023; Parmar et al., 2024; Zhou et al., 2024b; c; Zeng et al., 2024; Peng et al., 2024b; a; c; Gao et al., 2024; Zhu et al., 2024). In this study, we focus on real image editing, which allows users to freely modify actual photographs, producing highly realistic results. Typically, the inputs for image editing include an image and various conditions that help users accurately describe their desired changes. These conditions can encompass text prompts using natural language to specify the edits (Brooks et al., 2023; Zhang et al., 2024c; Fu et al., 2023; Zhang et al., 2024b), region masks to designate areas for modification (Zhao et al., 2024c; Zhuang et al., 2023; Wang et al., 2023), additional images to provide desired styles or objects (Chen et al., 2024; Lu et al., 2023; Yang et al., 2023), and drag points (Pan et al., 2023; Shi et al., 2024; Mou et al., 2023) that enable users to interactively move specific points in the image to target positions.
@@ Appendix B Watermarking Frequency Patterns
Figure 7: Frequency pattern visualizations for each watermarking method. Each subfigure displays the absolute difference between the Fourier spectrum $\mathcal{F}(\cdot)$ of the watermarked image $\bm{x}_{w}$ and the original image $\bm{x}_{o}$, i.e., $|\mathcal{F}(\bm{x}_{w})-\mathcal{F}(\bm{x}_{o})|$. The DWTDCT method is excluded because it closely resembles DWTDCTSVD and their pattern intensity is too weak to be discerned on the uniform scale. Please zoom in for a closer look.
@@
Figure 8: Impact of different editing methods on the frequency patterns of various watermarks. Each subfigure, analogous to Figure 3, displays the absolute difference between the Fourier spectrum $\mathcal{F}(\cdot)$ of the edited watermarked image $\epsilon(\bm{x}_{w})$ and the original image $\epsilon(\bm{x}_{o})$, i.e., $|\mathcal{F}(\epsilon(\bm{x}_{w}))-\mathcal{F}(\epsilon(\bm{x}_{o}))|$ to evaluate how the watermark patterns are altered by the editing process. The frequency patterns of VINE-R, VINE-B, MBRS, and StegaStamp are less affected compared to their original patterns (shown in Figure 7) than those of other watermarking methods. Please zoom in for a closer look.
Figure 8 demonstrates the effect of various editing methods on different watermark patterns within the frequency domain. As illustrated, the frequency patterns of VINE-R, VINE-B, MBRS, and StegaStamp are less affected compared to their original patterns (shown in Figure 7) than those of other watermarking methods. However, it challenging to isolate the effects on low-, mid-, and high-frequency bands when directly observing the watermarks.
Note that this finding is a byproduct of our design rather than a deliberate motivation of it. Our two key design elements (surrogate attack & generative prior adaptation) are based on the observations: (1) Image editing and image blurring exhibit similar frequency characteristics. (2) Image watermarking can be viewed as a form of conditional generation, where a generative prior can enhance image quality by making watermarks less visible. Regarding the intriguing finding—the robustness of our watermarking model against image editing being highly positively correlated with pattern intensity in the low-frequency region—this emerged from our training process rather than driving our design decisions. This finding aligns well with our results in Section 3.1, which show that image editing affects patterns in the high-frequency region more than those in the low-frequency region.
Appendix C Statistical Test
Let $\bm{w}\in\{0,1\}^{k}$ be the $k$-bit ground-truth watermark. The watermark $\bm{w}^{\prime}$ extracted from a watermarked image $\bm{x}_{w}$ is compared with the ground-truth $\bm{w}$ for detection. The detection statistical test relies on the number of matching bits $M(\bm{w},\bm{w}^{\prime})$: If
$M(\bm{w},\bm{w}^{\prime})\geq\tau,~{}~{}~{}~{}\text{where}~{}\tau\in\{0,1,2,\cdots,k\},$ (4)
then the image is flagged as watermarked. Formally, we test the statistical hypothesis $H_{1}$: ‘$\bm{x}$ contains the watermark $\bm{w}$’ against the null hypothesis $H_{0}$: ‘$\bm{x}$ does not contain the watermark $\bm{w}$’. Under $H_{0}$ (i.e., for original images), if the extracted bits $\bm{w}=\{w_{1}^{\prime},w_{2}^{\prime},\cdots,w_{k}^{\prime}\}$ (where $w_{i}^{\prime}$ is the i-th extracted bit) from a model are independent and identically distributed (i.i.d.) Bernoulli random variables with the matching probability $p_{o}$, then $M(\bm{w},\bm{w}^{\prime})$ follows a binomial distribution with parameters $(k,p_{o})$. This assumption is verified by Fernandez et al. (2023).
The false positive rate (FPR) is the probability that $M(\bm{w},\bm{w}^{\prime})$ takes a value bigger than the threshold $\tau$ under the null hypothesis $H_{0}$. It is obtained from the CDF of the binomial distribution, and a closed-form can be written with the regularized incomplete beta function $I_{p}(a,b)$ (Fernandez et al., 2023):
$\text{FPR}(\tau)=\mathbb{P}\left(M(\bm{w},\bm{w}^{\prime})>\tau|H_{0}\right)=\sum_{i=\tau+1}^{k}\binom{k}{i}p_{o}^{i}(1-p_{o})^{k-i}=I_{p_{o}}(\tau+1,k-\tau),$ (5)
where under $H_{0}$ (i.e., images without the watermark $\bm{w}$), $p_{o}$ should ideally be close to 0.5 to minimize the risk of false positive detection.
Similarly, the true positive rate (TPR) represents the probability that the number of matching bits exceeds the threshold $\tau$ under $H_{1}$, where the image contains the watermark. Thus, the TPR can be calculated by:
$\text{TPR}(\tau)=\mathbb{P}\left(M(\bm{w},\bm{w}^{\prime})>\tau|H_{1}\right)=\sum_{i=\tau+1}^{k}\binom{k}{i}p_{w}^{i}(1-p_{w})^{k-i}=I_{p_{w}}(\tau+1,k-\tau),$ (6)
where under $H_{1}$ (i.e., images with the watermark $\bm{w}$), $p_{w}$ should ideally be high enough (e.g., exceeding 0.8) to ensure the effectiveness of a watermarking model.
To further demonstrate that neither high bit accuracy nor AUROC alone guarantees a high TPR at a low FPR, consider the following example. Suppose we have a 100-bit watermarking model with a threshold $\tau$ of 70 to determine whether an image contains watermark $\bm{w}$. If the model extracts bits from watermarked images with a matching probability $p_{w}=0.8$ and from original images with a matching probability $p_{o}=0.5$, the resulting FPR would be $1.6\times 10^{-5}$ and the TPR would be 0.99. In this scenario, even though the bit accuracy for watermarked images is not exceptionally high (e.g., below 0.9), the model still achieves a high TPR at a very low FPR. In contrast, if another model has $p_{w}=0.9$ and $p_{o}=0.7$, achieving the same FPR would require setting the threshold $\tau$ to 87. Under these conditions, the TPR would only be 0.8. This example demonstrates that high bit accuracy for watermarked images does not necessarily ensure a high TPR when maintaining a low FPR. Therefore, relying solely on bit accuracy or AUROC may not be sufficient for achieving the desired performance in watermark detection.
Appendix D Resolution Scaling
D.1 Method
@@
D.2 Impact on watermarked image quality
@@
D.3 Impact on robustness against distortions
@@
Appendix E Condition Adapter Architecture
@@
Appendix F Implementation Details
F.1 Pre-training
@@ F.2 Fine-tuning
Appendix G Analysis of Image Editing
G.1 Rationale for Image Editing Selection
Image regeneration, global editing, local editing, and image-to-video generation cover the majority of editing needs. Image editing can be broadly divided into global and local categories, as user edits typically affect either the entire image or specific parts of it.
Global editing involves altering most of an image’s pixels while maintaining its overall layout, components, or semantics. Techniques such as style transfer, cartoonization, image translation, and scene transformation fall under this category and produce similar effects. For example, using prompts like ‘turn it into Van Gogh style’ or ‘convert it into a sketchy painting’ in an image editing model can effectively achieve style transfer.
Local editing, on the other hand, refers to modifications applied to specific elements, semantics, or regions within an image. This category includes image inpainting, image composition, object manipulation, attribute manipulation, and so forth.
While image regeneration and image-to-video generation are not strictly considered forms of image editing, they can be used to create similar digital content while removing watermarks, thereby posing a threat to copyright protection. For this reason, we have included them in our benchmark.
G.2 Image regeneration
@@ G.3 Global editing and local editing
Table 6: Comparison of watermarking performance in terms of detection accuracy on image-to-video generation with MAGE+. TPR@0.1%FPR is averaged over 1,000 videos.
Method TPR@0.1%FPR $\uparrow$ (%)
Frame 2 Frame 4 Frame 6 Frame 8 Frame 10 Average
MBRS (Jia et al., 2021) 89.57 88.67 87.45 86.51 84.42 87.32
CIN (Ma et al., 2022) 45.92 44.78 43.21 42.17 40.71 43.36
PIMoG (Fang et al., 2022) 78.23 76.99 75.72 74.91 73.12 75.79
RivaGAN (Zhang et al., 2019) 56.87 54.83 53.21 52.14 51.01 53.61
SepMark (Wu et al., 2023) 63.45 62.15 61.03 60.89 59.24 61.35
DWTDCT (Al-Haj, 2007) 30.57 29.51 28.89 27.72 26.87 28.71
DWTDCTSVD (Navas et al., 2008) 38.56 38.54 37.12 36.81 35.74 37.35
SSL (Fernandez et al., 2022) 81.21 80.95 78.99 77.18 76.12 78.89
StegaStamp (Tancik et al., 2020) 91.25 90.34 89.12 88.67 87.33 89.34
TrustMark (Bui et al., 2023) 90.35 90.12 89.45 87.69 86.13 88.75
EditGuard (Zhang et al., 2024d) 42.57 41.46 40.55 39.91 38.17 40.53
VINE-Base 92.22 91.35 90.74 89.12 88.01 90.29
VINE-Robust 93.14 92.88 91.32 90.27 89.12 91.35
G.4 Image-to-video generation
@@
Additionally, we add watermarks to 1,000 images from the CATER-GEN-v2 dataset Hu et al. (2022) and use the MAGE+ model Hu et al. (2022) to perform text-image-to-video (TI2V) generation, producing 10-frame videos. For watermark detection, we analyze frames 2, 4, 6, 8, and 10. The average detection accuracies across 1,000 videos are presented in Table 6.
Appendix H Inference Speed and GPU Memory Evaluation
@@
Appendix I Qualitative Comparison
@@
Appendix J Additional Benchmarking Results
@@
Figure 17: Qualitative comparison of the evaluated watermarking methods. The ‘Enlarged from Watermark’ column displays a magnified $40\times 40$ central region of the watermarked images, while the ‘Enlarged from Residual’ column shows a magnified $40\times 40$ central region of the residual images. Please zoom in for a closer examination.